OpenStudy (anonymous):

x/3 * 15/x+6. I need help understanding how to simply the expression.

OpenStudy (jdoe0001):

\(\bf \cfrac{x}{3}\times \cfrac{15}{x+6}\implies \cfrac{x\cdot 15}{3\cdot (x+6)}\)

OpenStudy (jdoe0001):

\(\bf \cfrac{x}{3}\times \cfrac{15}{x+6}\implies \cfrac{x\cdot 15}{3\cdot (x+6)}\implies \cfrac{15x}{3(x+6)}\implies \cfrac{(\cancel{3}\cdot 5)x}{\cancel{3}(x+6)}\)

OpenStudy (anonymous):

Thank you SO much. It totally makes sense :)